If North American casino roulette were fair, how much should it pay to a player who bets $100 on black and a black number comes up? The odds against winning a bet on a black in roulette are 1.01 to 1.

Answer by stanbon(75887)

If North American casino roulette were fair, how much should it pay to a player who bets $100 on black and a black number comes up? The odds against winning a bet on a black in roulette are 1.01 to 1.
----
With those odds P(losing) = 1.01/2.01
and P(winning) = 1/2.01
-----------------------------------------
Let "x" be payback to a player.
----
Note E(x) = zero if the game is "fair".
---
E(x) = (winnings)(prob of win) + (loss)(prob of loss)
E(x) = (x-100)(1/2.01) - 100(1.01/2.01) = 0
Multiply thru by 2.01 to get:
x-100 - 1.01*100 = 0
x - 100 - 101 = 0
x = $201
=================
Cheers,
Stan H.